A buffer consists of 0.27 M KHCO3 and 0.33 M K2CO3. Given that the K values for H2CO3 are, Ka1 = 4.5 x 10-7 and Ka2 = 4.7 x 10-11, calculate the pH for this buffer.
Answer #1

the pH will be simply given by the Henderson Hasselbalch equation :-

pH = pKa + log (conjugate base/acid)

pKa = - log (Ka1) = - log (4.5 x 10-7) = 6.34

acid = 0.27 M

base = 0.33 M

putting these values in the equation we get :-

pH = 6.34+ log (0.33/0.27)

pH = 6.34+0.087

pH = 6.43